OpenStudy (anonymous):

The linear density p in a rod 5 m long is 10/sqrt(x+4) kg/m, where x is measured in meters from one end of the rod. Find the average density of the rod. pave=? kg/m

OpenStudy (anonymous):

@dumbcow

OpenStudy (dumbcow):

you are given equation for density over an interval to find avg use this \[avg = \frac{1}{b-a} \int\limits_a ^b f(x) dx\] \[\frac{1}{5} \int\limits_0^5 \frac{10}{\sqrt{x+4}} dx\]

OpenStudy (anonymous):

so how does the P work into this?

OpenStudy (dumbcow):

not sure, i think its a label for density.... p = f(x) but it has no bearing on avg density

OpenStudy (anonymous):

I thought the question was saying it had to be plugged in somewhere. Thats what was throwing me haha

OpenStudy (dumbcow):

no the key is "x' is length from end of rod so you know you have to integrate wrt "x"

OpenStudy (dumbcow):

any problems integrating this? pretty straightforward with a "u" substitution

OpenStudy (anonymous):

oh ok i see where you are coming from

OpenStudy (anonymous):

naaa i got the integrating thing down now lol its just always the set up thats gets me

OpenStudy (dumbcow):

gotcha

OpenStudy (anonymous):

Thank you!

OpenStudy (dumbcow):

yw

OpenStudy (anonymous):

i got 4 as an answer

OpenStudy (dumbcow):

correct http://www.wolframalpha.com/input/?i=integrate+2%2Fsqrt%28x%2B4%29+from+0+to+5 useful tool for checking answers

OpenStudy (anonymous):

sweet I actually have another question just want you to tell me if im setting it up right find the average value of the function h on the interval. h(x)=6cos^4 x sin x [0.pi] the integral if just \[\frac{ 1 }{ \pi }\int\limits_{0}^{\pi}6\cos^4xsinx \]

OpenStudy (dumbcow):

looks good

OpenStudy (anonymous):

ok i think im starting to get this

OpenStudy (anonymous):

Now all i need to get is those cylindircal shells and i can take the test lol

OpenStudy (dumbcow):

oh yeah have you learned shell method now

OpenStudy (anonymous):

yep Use the method of cylindrical shells to find the volume of the solid obtained by rotatin the region bounded by the given curves about the x-axis. xy=5 x=0 y=5 y=7 so x=5/y \[\int\limits_{5}^{7}(2\pi y)(5/y) = 20\pi\] i think.... lol

OpenStudy (dumbcow):

hold on let me draw it out lol

OpenStudy (dumbcow):

ok you got it

OpenStudy (anonymous):

:D happy danceeee

OpenStudy (anonymous):

ok will you check one more for me? find the volume v of the described solid S. The base of S is a circular disk with radius 5r. Parallel cross sections perpendicular to he base are squares. i got V=2000/3 r^3

OpenStudy (dumbcow):

i believe you are right :) sum up all the areas....area of square is 4y^2...put y in terms of x ....

OpenStudy (anonymous):

Perfect :) you rock my friend

OpenStudy (dumbcow):

thank you...looks like you got the hang of it, keep practicing

OpenStudy (dumbcow):

here is link to good reference site with notes/examples/practice problems http://tutorial.math.lamar.edu/Classes/CalcII/CalcII.aspx

OpenStudy (anonymous):

ohhhh nice I hope it has something about J thats my new challenge lol